# A classification of Prüfer domains of integer-valued polynomials on algebras

### v1

@@ 4.1 Building a Prüfer domain inside $K[X]$
As before, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$. First we show that $g$ satisfies either (i) or (ii) for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$. Fix $P$ and $i\in I_{P}$, let $V=V_{P,i}$, $M=M_{P,i}$, and $v=v_{P,i}$. By Proposition 4.5(2), $f(f^{r-1}-1)=f^{r}-f\in\mathfrak{M}$ for each unitary maximal ideal of $\mathscr{D}$ above $\mathfrak{p}$. So, either $v(f)>0$ or $v(f^{r-1}-1)>0$. Note that these two conditions cannot occur simultaneously.
Assume first that $v(f)>0$. Then, $v(g)=v(\tfrac{f}{\mathfrak{p}})$. If (i) holds for $f$, then $v(f)=tv(\mathfrak{p})$ for some integer $t\geq 1$. So, $v(g)=(t-1)v(\mathfrak{p})$ and (i) holds for $g$. On the other hand, if $f$ satisfies (ii), then so does $g$, because for all $t\geq 0$ we have $v(\tfrac{g}{\mathfrak{p}^{t}})=v(\tfrac{f}{\mathfrak{p}^{t+1}})>0$.
Next, assume that $v(f^{r-1}-1)>0$. Then, $f^{r-1}-1\in M$, $(f^{r-1}-1)^{s}\in\mathfrak{p}V$ and $v(f)=0$ (so, in this case, condition (i) is true for $f$ with $t=0$). By Lemma 4.6, either condition (i) or condition (ii) holds for $(f^{r-1}-1)^{s}$. Moreover, if (i) is true for $(f^{r-1}-1)^{s}$, then the fact that $(f^{r-1}-1)^{s}\in\mathfrak{p}V$ means $v((f^{r-1}-1)^{s}/\mathfrak{p}^{t})=0$ for some $t\geq 1$. We have $v(g)=v((f^{r-1}-1)^{s}/\mathfrak{p})$, so either (i) or (ii) holds for $g$.
Finally, the fact that $g$ satisfies either (i) or (ii) implies that $g\in V$, and this is true for every $P|\mathfrak{p}D_{0}$ and every $i\in I_{P}$. This is enough to conclude that $g\in\mathscr{D}$, because if $P^{\prime}$ is a nonzero prime of $D$ that is not over $\mathfrak{p}$, then $\mathfrak{p}$ is a unit in each valuation overring associated to $P^{\prime}$. ∎
@@
Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $\mathfrak{p}$ and let $f\in\mathscr{D}\setminus\{0\}$. Then, either $f^{s_{\mathfrak{p}}}/\mathfrak{p}^{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all integers $t\geq 1$, or $f^{s_{\mathfrak{p}}}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$ for some integer $t\geq 0$.
@@
Once again, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$. There is nothing to prove if $f\notin\mathfrak{M}$, so assume that $f\in\mathfrak{M}$. Let $f_{0}=f^{s}$, and for each $k\geq 1$, let
$f_{k}=\dfrac{f_{k-1}(f_{k-1}^{r-1}-1)^{s}}{\mathfrak{p}}.$
@@
For each $k\geq 1$, let $h_{k}=\prod_{i=0}^{k-1}(f_{i}^{r-1}-1)^{s}$. Then, $f_{k}=(f^{s}h_{k})/\mathfrak{p}^{k}$ for all $k\geq 1$. Notice that if $f_{i}\in\mathfrak{M}$ for all $0\leq i\leq k-1$, then $h_{k}\notin\mathfrak{M}$. Hence, if there exists a smallest positive integer $t$ such that $f_{t}\notin\mathfrak{M}$, then $f^{s}/\mathfrak{p}^{t}=f_{t}/h_{t}$ is a unit in $\mathscr{D}_{\mathfrak{M}}$. Otherwise, $f_{t}\in\mathfrak{M}$ for every $t\geq 1$, each $h_{t}\notin\mathfrak{M}$, and $f^{s}/\mathfrak{p}^{t}=f_{t}/h_{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all $t\geq 1$. ∎
@@
Let $\mathfrak{M}$ be a maximal ideal of $\mathscr{D}$. If $\mathfrak{M}$ is non-unitary, then $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain by Lemma 4.3. So, assume that $\mathfrak{M}$ is a unitary maximal ideal above $\mathfrak{p}$. Let $f,g\in\mathscr{D}$ be nonzero and coprime in $K[X]$. We will show that either $f/g$ or $g/f$ is in $\mathscr{D}_{\mathfrak{M}}$. This is clear if either $f\notin\mathfrak{M}$ or $g\notin\mathfrak{M}$. So, assume that both $f$ and $g$ are in $\mathfrak{M}$.
Let $s=s_{\mathfrak{p}}$. Since $f$ and $g$ are coprime in $K[X]$, there exist $h_{1},h_{2}\in K[X]$ such that $1=h_{1}f^{s}+h_{2}g^{s}$. Choose $d\in D\setminus\{0\}$ such that $dh_{1},dh_{2}\in D[X]\subseteq\mathscr{D}$. If both $f^{s}/\mathfrak{p}^{t}$ and $g^{s}/\mathfrak{p}^{t}$ are in $\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for every positive integer $t$, then
$\dfrac{d}{\mathfrak{p}^{t}}=\dfrac{dh_{1}f^{s}+dh_{2}g^{s}}{\mathfrak{p}^{t}}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$
for all $t\geq 1$. This implies that $d/\mathfrak{p}^{t}\in P=\mathfrak{M}\cap D$ for all $t\geq 1$, which is impossible. By Lemma 4.8, we can find $t\geq 1$ such that both $f^{s}/\mathfrak{p}^{t}$ and $g^{s}/\mathfrak{p}^{t}$ are in $\mathscr{D}_{\mathfrak{M}}$, and either $f^{s}/\mathfrak{p}^{t}$ or $g^{s}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$. Without loss of generality, assume that $f^{s}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$. Then,
$(g/f)^{s}=g^{s}/f^{s}=(g^{s}/\mathfrak{p}^{t})/(f^{s}/\mathfrak{p}^{t})\in\mathscr{D}_{\mathfrak{M}}.$ (4.10)
@@ 4.2 Proving that $\textnormal{Int}_{K}(\Lambda_{n})$ is Prufer
(1) Let $n\geq 1$ and express $\textnormal{Int}_{K}(\Lambda_{n})$ as in Theorem 4.11. Let $\mathfrak{p}$ be a nonzero prime element of $D_{0}$. Since $D$ is an ADDB-domain, we know that the integers $E_{\mathfrak{p}}$ and $F_{\mathfrak{p}}$ from Notation 1.1 are bounded above, say by some integer $d_{\mathfrak{p}}$. Let $P$ be a prime of $D$ above $\mathfrak{p}$. By the results of [24, Proposition 2.2], we have $M_{P,\alpha}^{n}=PV_{P,\alpha}$ and $V_{P,\alpha}/M_{P,\alpha}:D/P]\leq n$ for each $\alpha\in\widehat{\Lambda}_{P,n}$. Therefore, $\textnormal{Int}_{K}(\Lambda_{n})$ is a domain formed by Construction 4.1, because we can take $e_{\mathfrak{p}}=f_{\mathfrak{p}}=nd_{\mathfrak{p}}$. Hence, by Theorem 4.9, $\textnormal{Int}_{K}(\Lambda_{n})$ is a Prüfer domain.
@@
We prove that $\textnormal{Int}_{K}(A,A^{\prime})_{P}\subseteq\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$ by employing an induction technique due to Rush [32, Proposition 1.4]. The proof that $\textnormal{Int}_{K}(A)_{P}\subseteq\textnormal{Int}_{K}(A_{P})$ is similar. Let $f\in\textnormal{Int}_{K}(A,A^{\prime})_{P}$ have degree $m$ and let $a/d\in A_{P}$, where $a\in A$ and $d\in D\setminus P$. We use induction on $m$. If $m=0$, then $f\in D_{P}\subseteq A_{P}^{\prime}$ and we are done. Assume that $m\geq 1$ and that any polynomial in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$ of degree at most m1m-1 is an element of $\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$. Let $g(x)=d^{m}f(x)-f(dx)$. Then, both $f(dx)$ and $g$ are in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$, and $\deg g\leq m-1$. By induction, $g\in\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$.

### v2

@@ 4.1 Building a Prüfer domain inside $K[X]$
As before, let $r=r_{p}$ and $s=s_{p}$. First we show that $g$ satisfies either (i) or (ii) for each $P|pD_{0}$ and each $i\in I_{P}$. Fix $P$ and $i\in I_{P}$, let $V=V_{P,i}$, $M=M_{P,i}$, and $v=v_{P,i}$. By Proposition 4.5(2), $f(f^{r-1}-1)=f^{r}-f\in\mathfrak{M}$ for each unitary maximal ideal of $\mathscr{D}$ above $p$. So, either $v(f)>0$ or $v(f^{r-1}-1)>0$. Note that these two conditions cannot occur simultaneously.
Assume first that $v(f)>0$. Then, $v(g)=v(\tfrac{f}{p})$. If (i) holds for $f$, then $v(f)=tv(p)$ for some integer $t\geq 1$. So, v(g)=(t1)v(p)v(g)=(t-1)v(p) and (i) holds for $g$. On the other hand, if $f$ satisfies (ii), then so does $g$, because for all $t\geq 0$ we have $v(\tfrac{g}{p^{t}})=v(\tfrac{f}{p^{t+1}})>0$.
Next, assume that $v(f^{r-1}-1)>0$. Then, $f^{r-1}-1\in M$, $(f^{r-1}-1)^{s}\in pV$ and $v(f)=0$ (so, in this case, condition (i) is true for $f$ with $t=0$). By Lemma 4.6, either condition (i) or condition (ii) holds for $(f^{r-1}-1)^{s}$. Moreover, if (i) is true for $(f^{r-1}-1)^{s}$, then the fact that $(f^{r-1}-1)^{s}\in pV$ means $v((f^{r-1}-1)^{s}/p^{t})=0$ for some $t\geq 1$. We have $v(g)=v((f^{r-1}-1)^{s}/p)$, so either (i) or (ii) holds for $g$.
Finally, the fact that $g$ satisfies either (i) or (ii) implies that $g\in V$, and this is true for every $P|pD_{0}$ and every $i\in I_{P}$. This is enough to conclude that $g\in\mathscr{D}$, because if $P^{\prime}$ is a nonzero prime of $D$ that is not over $p$, then $p$ is a unit in each valuation overring associated to $P^{\prime}$. ∎
@@
Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $p$ and let $f\in\mathscr{D}\setminus\{0\}$. Then, either $f^{s_{p}}/p^{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all integers $t\geq 1$, or $f^{s_{p}}/p^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$ for some integer $t\geq 0$.
@@
Once again, let $r=r_{p}$ and $s=s_{p}$. There is nothing to prove if $f\notin\mathfrak{M}$, so assume that $f\in\mathfrak{M}$. Let $f_{0}=f^{s}$, and for each $k\geq 1$, let
$f_{k}=\dfrac{f_{k-1}(f_{k-1}^{r-1}-1)^{s}}{p}.$
@@
For each $k\geq 1$, let $h_{k}=\prod_{i=0}^{k-1}(f_{i}^{r-1}-1)^{s}$. Then, $f_{k}=(f^{s}h_{k})/p^{k}$ for all $k\geq 1$. Notice that if $f_{i}\in\mathfrak{M}$ for all $0\leq i\leq k-1$, then $h_{k}\notin\mathfrak{M}$. Hence, if there exists a smallest positive integer $t$ such that $f_{t}\notin\mathfrak{M}$, then $f^{s}/p^{t}=f_{t}/h_{t}$ is a unit in $\mathscr{D}_{\mathfrak{M}}$. Otherwise, $f_{t}\in\mathfrak{M}$ for every $t\geq 1$, each $h_{t}\notin\mathfrak{M}$, and $f^{s}/p^{t}=f_{t}/h_{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all $t\geq 1$. ∎
@@
Let $\mathfrak{M}$ be a maximal ideal of $\mathscr{D}$. If $\mathfrak{M}$ is non-unitary, then $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain by Lemma 4.3. So, assume that $\mathfrak{M}$ is a unitary maximal ideal above $p$. Let $f,g\in\mathscr{D}$ be nonzero and coprime in $K[X]$. We will show that either $f/g$ or $g/f$ is in $\mathscr{D}_{\mathfrak{M}}$. This is clear if either $f\notin\mathfrak{M}$ or $g\notin\mathfrak{M}$. So, assume that both $f$ and $g$ are in $\mathfrak{M}$.
Let $s=s_{p}$. Since $f$ and $g$ are coprime in $K[X]$, there exist $h_{1},h_{2}\in K[X]$ such that $1=h_{1}f^{s}+h_{2}g^{s}$. Choose $d\in D\setminus\{0\}$ such that $dh_{1},dh_{2}\in D[X]\subseteq\mathscr{D}$. If both $f^{s}/p^{t}$ and $g^{s}/p^{t}$ are in $\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for every positive integer $t$, then
$\dfrac{d}{p^{t}}=\dfrac{dh_{1}f^{s}+dh_{2}g^{s}}{p^{t}}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$
for all $t\geq 1$. This implies that $d/p^{t}\in P=\mathfrak{M}\cap D$ for all $t\geq 1$, which is impossible. By Lemma 4.8, we can find $t\geq 1$ such that both $f^{s}/p^{t}$ and $g^{s}/p^{t}$ are in $\mathscr{D}_{\mathfrak{M}}$, and either $f^{s}/p^{t}$ or $g^{s}/p^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$. Without loss of generality, assume that $f^{s}/p^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$. Then,
$(g/f)^{s}=g^{s}/f^{s}=(g^{s}/p^{t})/(f^{s}/p^{t})\in\mathscr{D}_{\mathfrak{M}}.$ (4.10)
@@ 4.2 Proving that $\textnormal{Int}_{K}(\Lambda_{n})$ is Prufer
(1) Let $n\geq 1$ and express $\textnormal{Int}_{K}(\Lambda_{n})$ as in Theorem 4.11. Let $p$ be a nonzero prime element of $D_{0}$. Since $D$ is an ADDB-domain, we know that the integers $E_{p}$ and $F_{p}$ from Notation 1.1 are bounded above, say by some integer $d_{p}$. Let $P$ be a prime of $D$ above $p$. By the results of [24, Proposition 2.2], we have $M_{P,\alpha}^{e_{P,\alpha}}=PV_{P,\alpha}$ for some $e_{P,\alpha}\leq n$ and $V_{P,\alpha}/M_{P,\alpha}:D/P]\leq n$ for each $\alpha\in\widehat{\Lambda}_{P,n}$. Therefore, $\textnormal{Int}_{K}(\Lambda_{n})$ is a domain formed by Construction 4.1, because we can take $e_{p}=f_{p}=nd_{p}$. Hence, by Theorem 4.9, $\textnormal{Int}_{K}(\Lambda_{n})$ is a Prüfer domain.
@@
We prove that $\textnormal{Int}_{K}(A,A^{\prime})_{P}\subseteq\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$ by employing an induction technique due to Rush [32, Proposition 1.4]. The proof that $\textnormal{Int}_{K}(A)_{P}\subseteq\textnormal{Int}_{K}(A_{P})$ is similar. Let $f\in\textnormal{Int}_{K}(A,A^{\prime})_{P}$ have degree $m$. We use induction on $m$. If $m=0$, then $f\in A_{P}^{\prime}$ and we are done. Assume that $m\geq 1$ and that any polynomial in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$ of degree at most m1m-1 is an element of $\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$. Let $a/d\in A_{P}$, where $a\in A$ and $d\in D\setminus P$. Let $g(x)=d^{m}f(x)-f(dx)$. Then, both $f(dx)$ and $g$ are in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$, and $\deg g\leq m-1$. By induction, $g\in\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$.